If Mary can address a box of envelopes in 5 hours and jane can address in the same box of envelopes in 10 hours, how many minutes will it take Mary and Jane working together to address all the envelopes in the box?
Answer by richwmiller(17219)

1/5+1/10=1/x
2/10+1/10=1/x
3/10=1/x
x=10/3